Suppose you bought 500 shares of stock at an initial price of $44 per share. The stock paid a dividend of $.42 per share during the following year, and the share price at the end of the year was $45.

a.

What is the capital gains yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Omit the "%" sign in your response.)

  Capital gains yield %
b.

What is the dividend yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Omit the "%" sign in your response.)

  Dividend yield %
c.

What is the total rate of return on the investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Omit the"%"sign in your response.)

  Total rate of return %
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars
a)
Capital Gain Yield = (P1-P0)/P0
Capital Gain Yield = (45 - 44)/44
Capital Gain Yield = 2.27%
b)
Dividend Yield = Most recent dividend/Current Share Price
Dividend Yield = .42/45
Dividend Yield = .93%
c)
Rate of return = (1 + .42)/44 = 3.23%